OpenStudy (anonymous):

What is the solution to 3 times the square root of 4 minus 2 times the square root of 4?

OpenStudy (mrnood):

\[(3\times \sqrt{4})-(2\times \sqrt{4})\] First What is square root of 4?

OpenStudy (mrnood):

Then - solve the above equation